A mathematical biologist created a model to administer medicine reaction (measured in change of blood pressure or temperature) with the model given by 

2

R= m^2 (C/2- M/3)

where c is a positive constant and m is the amount of medicine absorbed into the blood. The sensitivity to the medication is defined to be the rate of change of reaction R with respect to the amount of medicine absorbed in the blood.


1
Expert's answer
2021-11-08T20:44:55-0500
"R=m^2\big(\dfrac{c}{2}-\dfrac{m}{3}\big)"

(a)


"R'=\big(m^2\big(\dfrac{c}{2}-\dfrac{m}{3}\big)\big)'=cm-m^2"

(b)


"R'=0=>cm-m^2=0"

"m=0\ or\ m=c, c>0"

If "0<m<c," then "R'>0, R" increases.

If "m>c," then "R'<0, R" decreases.

The reaction is maximum at "m=c"


"R(c)=c^2\big(\dfrac{c}{2}-\dfrac{c}{3}\big)=\dfrac{c^3}{6}"

(c)


"(R')'=(cm-m^2)'=c-2m"

The instantaneous rate of change of sensitivity is the second derivative of the reaction


"(R')'=c-2m"

(d)


"(R')'=0=>c-2m=0=>m=\dfrac{c}{2}"

If "0<m<\dfrac{c}{2}," then "(R')'>0, R'" increases.

If "m>\dfrac{c}{2}," then "(R')'<0, R'" decreases.


The sensitivity is maximum at "m=\dfrac{c}{2}"


"R'(\dfrac{c}{2})=c(\dfrac{c}{2})-(\dfrac{c}{2})^2=\dfrac{c^2}{4}"